OpenStudy (anonymous):

Solve the following Quadratic Equations? 2x2 – 15x – 8 = 0

OpenStudy (anonymous):

2x^2 - 15x - 8 = 0. (2x + 1)(x - 8) = 0. 2x + 1 = 0. 2x = - 1. x = - 1/2. One Answer. x - 8 = 0. x = 8. Second Answer.

OpenStudy (callisto):

2x^2 – 15x – 8 = 0 2x^2 -16x + x - 8 = 0 2x(x-8) + (x-8) = 0 (2x+1) ( x-8) = 0 So, 2x+1 = 0 or x-8 = 0 <- solve these two equations.

OpenStudy (callisto):

@ashna Don't provide answer without interaction with the asker or do all the work for him/her. Thanks!

OpenStudy (anonymous):

he is doing well i better understand his way!!

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

@Callisto i better understand his way!!! THANKKK YOU!!!

OpenStudy (callisto):

@broderick365 We are probably doing the same thing, but he/she provides the final answer as well.